OpenStudy (anonymous):

Solve x = log9^4. Can someone explain how to solve these things? :/ X_X

OpenStudy (anonymous):

\[loga^m = mloga \Rightarrow \log9^4 = 4\log9\]

OpenStudy (johnweldon1993):

exactly as @sachinrajsharma said :)

OpenStudy (anonymous):

but I'm supposed to find x, how to do that? :/

OpenStudy (johnweldon1993):

that would be what 4log9 equates to....ill do the calculation for you 4log9 = 3.816970038 Thats what x equals :)

OpenStudy (anonymous):

x = 4log9 = 4log3^2 = 8log3 now put the value of log3 u get ur answer

OpenStudy (anonymous):

Thanks for the answer but I'd like to understand how to solve them, I still don't get it X_X How did you get to 8log3?

OpenStudy (johnweldon1993):

remember the equation @sachinrajsharma put up there? \[\log_{a}^m = m \log_{}a \] That means when it goes from x = 4log9 to x = 4log3^2 (3^2 is just another way of writing 9 right?) the equation says you multiply the exponent (2) to the log 4log3.....now since you already have 4 as a coefficent....you multiply 4 and 2...to get 8...which leaves you with 8log3

OpenStudy (johnweldon1993):

equation didn't come out right...but it's the equation they put up there

OpenStudy (anonymous):

Ohhhh ok I see. Sorry I'm kinda slow. And I'm still not sure I understand exactly why x is that number but thanks for explaining.

OpenStudy (johnweldon1993):

you're not slow...i'm sure you're very intelligent :) and x equals that because x = 4log9 it just say x is equal to whatever 4log9 equates to... just like x = 2...that says x = 2 x = 6(10) means x = 60.... so x is whatever 4log9 equals :)

OpenStudy (anonymous):

Oh damn. I misunderstood everything about log I guess lol Get it now :) ty

OpenStudy (johnweldon1993):

haha good :) and anytime :)